A man weighing 660 N and a woman weighing 460 N have the same momentum. What is the ratio of the man's kinetic energy Km to that of the woman Kw?
Answer #1

here,

mass of man , m1 = weight of man /g

m1 = 660 /9.81 kg = 67.28 kg

mass of Woman , m2 = weight of Woman /g

m2 = 460 /9.81 kg = 46.89 kg

let the velocity of man and woman be v1 and v2

as they have same momentum

m1 * v1 = m2 * v2

v1 /v2 = m2 /m1

v1 /v2 = 46.89 /67.28

v1 /v2 = 0.697

the ratio of the man's kinetic energy Km to that of the woman Kw , R = Km /Kw

R = (0.5 * m1 * v1^2) /(0.5 * m2 * v2^2)

R = (67.28/46.89) * ( 0.697)^2

R = 0.7

the ratio of the man's kinetic energy Km to that of the woman Kw is 0.7